OpenStudy (anonymous):

This table shows a 2 day race. Day 1: 7:57:54 Day 2: 6:59:56 What was the total time for the two day race? The answer is 14:57:50, but I don't understand how to add time? Thanks :)

OpenStudy (jack1):

Day 1: 7:57:54 Day 2: 6:59:56 add, round off at 60 56+54 = 1:50 57+59 + 1= 1:57 7+6 +1 = 14 14 hrs 57 mins 50 secs

OpenStudy (anonymous):

Thanks Jack1, makes so much more sense now :)